# Design of superparamagnetic nanoparticle-materials for high-frequency inductor cores

### v3

@@ III Evaluation of particles for nanocomposites
Table 1 lists calculated susceptibilities and losses at 2 MHz for composites of 30 volume % 201 nm FeNi3 particles with anisotropy axes oriented parallel, perpendicularly and randomly wrt. the applied field. The losses are calculated with an applied field amplitude such that a flux density amplitude of 30 mT is achieved in the nanoparticle composites. From the calculated losses in Table 1, it is seen that for both the parallel and perpendicularly aligned cases, losses lower than 200 mW/cm3 are possible. The aligned particles have lower losses at a given flux density than the random case due to lower applied field amplitude used to reach the desired flux density. The effect of temperature on the power loss at a given flux density can be calculated from the derived theory above. Susceptibility will in general decrease as , while the out-of-phase component will decrease further due to lower and . Thus higher field will be needed for higher temperature to reach desired flux density, but the power loss will decrease. Based on these calculations, the susceptibility and losses of perpendicularly and parallel aligned, spherical single-domain FeNi3 nanoparticles seem promising for micro-inductor applications.
The direct effect of dipolar interaction on the susceptibility can be illustrated by use of equation (14). We find that for a hypothetical FeNi3 201 nm diameter two-particle system, an increase (/decrease) of susceptibility due to the two particles being oriented along a chain parallel (/perpendicular) to the applied field will be of the order of 0.2 (0.1) % of the total susceptibility at 30 vol% (assuming that the volume fraction is ). The effect is seen to be rather small, and the change holds true for particles with similar high moment. Hence, we deem direct interaction effects on susceptibility to be of low importance for materials with 30 vol% of particles, as long as the particles are still superparamagnetic, see below.
The effect of dipolar interactions on the superparamagnetic relaxation is not yet fully described in the literature [38]. It has been suggested that dipolar interactions can slightly lower the effective anisotropy barrier in case of low volume fraction [39], while for dense aggregates it is often said to drastically increase effective anisotropy, as it corroborates with suppression of superparamagnetism [19, 40, 41, 38]. From an energy-consideration one could argue that it is favourable to have larger particle size with lower volume fraction, as the increase in uniaxial anisotropy energy is lower than the decrease in dipolar interaction energy, at a fixed nanocomposite susceptibility.
@@ IV Discussion
In this work, we have shown how to calculate the susceptibility for magnetic nanoparticle materials. The presented derivations show results for particle alignment, cubic anisotropy, and effects from dipolar interactions not given before [42, 43]. Moreover, the found dependency of particle susceptibility on the applied field direction wrt. the easy axis for uniaxial anisotropy case is interesting for application within power electronics magnetics, as it can be used to tune susceptibility and losses for different materials. We note that the susceptibility of cubic anisotropy particles cannot be tuned by alignment, and since other magnetic properties, like coercive field, are rather similar for cubic and random uniaxial case [44], it will only be possible to study these effects when particles are aligned [25, 13].
@@
Magnetic nanoparticles have previously been investigated experimentally for use as inductor core materials in the MHz range [7, 8, 9, 10, 11, 12, 13, 14, 18, 16], cf. Fig. 1. By use of our theoretical framework, we are able to predict susceptibility values quantitatively and in general trend with the experimental reports. Indeed, given the uncertainties in particle size distributions and material parameters (especially anisotropy and saturation magnetization, which may vary with size), some variation exist between the values from the model and the experiments, but overall there is a strong match. This supports the predictive power of our model.
@@
When comparing the experimental in-phase susceptibilities, shown in Fig. 1, with our model, we find that most studies so far have not optimised particle materials for the use case. Hence, there is room for improvements, i.e., to reach into the region of interest shown in turquoise in Fig. 1. For instance, for most materials, the nanoparticles were packed densely or were aggregated, which effectively renders the particles blocked and results in large coercive field of several thousand kA/m [7, 10, 13, 14]. Consequently, the resulting materials behave much like if one uses larger particles, i.e. with lower susceptibility and larger coercivity, see for example [11]. Some materials show no or low hysteresis [8, 9], which should be a sign of superparamagnetism, but these materials have only relatively low susceptibilities. This seems to be due to too low saturation magnetisation of the used particles and/or too dilute particle concentrations. One reference reports a high particle susceptibility of 122, using relatively densely packed 8 nm Fe particles [12], which fits well with the results from Fig. 4, but with seemingly lower particle anisotropy than that we have used for the calculations.
@@
Nanocomposite materials thus show the potential to reach the area of interest of susceptibility above 50 in the frequency range above 2 MHz as indicated on Fig. 1. Based on our model, we propose that optimised particles, i.e. spherical 8 nm Fe (or FeCo) particles or 20 nm FeNi3 particles, have the required properties for inductor core applications. A limitation of the derived theoretic expressions is, that effective particle property values are somewhat sparse. While we have used reported effective particle values in our evaluation of superparamagnetic particles, it is not necessary possible to make particles in the desired size range (optimally monodisperse) with the given properties or orientation wrt. the applied field.
@@
Alignment of particles with the effective easy axis parallel to the applied field has been investigated experimentally [25, 13], but it is not clear if the found increase in susceptibility could be due to aggregation or chain formation. Any aggregation or chain formation might render the particles blocked or change their dynamic behaviour [38]. Magnetic interactions in derivation of susceptibility have, to our knowledge, only been introduced as effective field [43], where susceptibility has been found to increase. Monte-Carlo simulations have shown a decrease in susceptibility for increasing volume fraction of particles [45]. We find, that for superparamagnetic particles the interaction depend on particle-assembly orientation wrt. the applied field, as shown in equation (14) and therefore depends on particle packing.
@@
An relevant study case would be an uniaxial anisotropy particle with an added shape anisotropy axis perpendicular to the applied field direction. This combination might have an lower effective anisotropy barrier than the pure uniaxial anisotropy case. Another interesting case would be prolate particles, as there for such would be an energy minimum laying at a plane, such that almost no barrier needs to be overcome to reverse the magnetisation in the plane of the oblate particle [46].polar interactions changes the susceptibility, and could potentially also change the frequency dependence due to increased effective anisotropy. We note is that many materials presented in Fig. 1 are relatively dense materials, and reported susceptibilities of low coercivity materials still seem to have stable susceptibility, although low, even at high frequencies [9, 12]. It has been shown that the applied field can change the superparamagnetic relaxation, allowing for faster response in applied field than without [47, 31]. This could explain the high blocking frequency seen for densely packed materials [12, 9]
@@ V CONCLUSIONS
We have developed a comprehensive statistical mechanics method to calculate nanoparticle susceptibility including effects of size, shape, anisotropy, saturation magnetization, and inter-particle interactions.
Using the derived theory, we find that nanoparticles with large saturation magnetisation and low anisotropy are the most suited for inductor cores in power electronics applications. Moreover, we show that the particles should have an optimised diameter, such that they are as large as possible but remain superparamagnetic at the operation frequency. For this, a narrow particle size distribution is preferable. In relation to particle shape, we find that elongation can increase susceptibility to some degree, but due to the increase in effective anisotropy, it is accompanied by a large increase in out-of-phase component (i.e. magnetic hysteresis losses). Therefore spherical or near-spherical particle shapes would be best for inductor cores. er induced magnetisation have been found to be favourable for materials where the uniaxial anisotropy axes of the particles are aligned compared to the random case.
As an example of optimised particles, we show that a material with 30 vol% of spherical 201 nm FeNi3 particles could potentially have a susceptibility above 209 (487 for aligned particles) while lower losses than state-of-the-art ferrite magnetic core materials at 2 MHz operation.
Comparing the findings from our theoretical model to reported experimental results, we find there is good agreement, confirming the predictive power of the model. Lower susceptibilities (´20-50), compared to the optimal cases for inductors, has been reported for nanocomposites in most experimental studies. This appear to be due to non-optimized usage of low magnetisation materials, too large/too small particles, too high/too low volume fraction of particles, or too dense aggregates with dipolar interactions, which can diminish the susceptibility. Hence better composite materials seems possible with interesting potential for use in micro-fabricated inductor cores at high frequencies where the bulk ferrite materials like TDK’s PC200 cannot operate.
@@ Acknowledgements.
The authors thank Ron B. Goldfarb for stimulating discussions, and the Independent Research Fund Denmark for financial support (project HiFMag, grant number 9041-00231A).References
- Yawger [2022] R. Yawger, Psma power technology roadmap [psma corner], Ieee Power Electronics Magazine 9, 10 (2022).
@@
- Stoner and Wohlfarth [1948] E. Stoner and E. Wohlfarth, A mechanism of magnetic hysteresis in heterogeneous alloys, Philosophical Transactions of the Royal Society of London Series A-mathematical and Physical Sciences 240, 599 (1948).
@@
- Walker et al. [1993] M. Walker, P. Mayo, K. OGrady, S. Charles, and R. Chantrell, The magnetic-properties of single-domain particles with cubic anisotropy .1. hysteresis loops, Journal of Physics-condensed Matter 5, 2779 (1993).
@@
- Ludwig et al. [2017] F. Ludwig, C. Balceris, and C. Johansson, The anisotropy of the ac susceptibility of immobilized magnetic nanoparticles-the influence of intra-potential-well contribution on the ac susceptibility spectrum, Ieee Transactions on Magnetics 53, 7894242 (2017).
- Carrey et al. [2011] J. Carrey, B. Mehdaoui, and M. Respaud, Erratum: Simple models for dynamic hysteresis loop calculations of magnetic single-domain nanoparticles: Application to magnetic hyperthermia optimization (journal of applied physics (2011) 109 (083921)), Journal of Applied Physics 110, 039902 (2011).
- Raikher and Shliomis [1974] Y. L. Raikher and M. I. Shliomis, Dispersion theory of the magnetic susceptibility of small ferromagnetic particles, Zhurnal Eksperimental’noi I Teoreticheskoi Fiziki 67, 1060 (1974).
- Shliomis and Stepanov [1993] M. Shliomis and V. Stepanov, Frequency-dependence and long-time relaxation of the susceptibility of the magnetic fluids, Journal of Magnetism and Magnetic Materials 122, 176 (1993).
- EISENSTEIN and AHARONI [1977] I. EISENSTEIN and A. AHARONI, High-order superparamagnetic relaxation-times, Physica B and C 86, 1429 (1977).
@@
- Bohra et al. [2021] M. Bohra, S. V. Battula, V. Alman, A. Annadi, and V. Singh, Design of various ni–cr nanostructures and deducing their magnetic anisotropy, Applied Nanoscience (switzerland) , 1 (2021).
@@
- Kumari et al. [2021] K. Kumari, A. Kumar, J. E. Lee, and B. H. Koo, Investigating the origin of exchange bias effect in ferromagnetic feni nanoparticles prepared via controlled synthesis, Applied Nanoscience (switzerland) , 1 (2021).
- Durhuus et al. [2025] F. L. Durhuus, T. van Biljvelt Rix, M. A. Glod, M. Beleggia, and C. Frandsen, Neel relaxation of magnetic nanoparticle clusters, IEEE Magnetics Letters (accepted) (2025).

### v4

@@ III Particle optimization for nanocomposites
Table 1 lists calculated susceptibilities and losses at 2 MHz for nanocomposites of 30 volume % 201 nm FeNi3 particles with anisotropy axes oriented parallel, perpendicular and randomly with respect to the applied field. The losses are calculated with an applied field amplitude such that a flux density amplitude of 30 mT is achieved in the nanocomposites. The calculated losses in Table 1 indicate that for both the parallel and perpendicularly aligned cases, losses lower than 200 mW/cm3 are possible. The aligned particles have lower losses at a given flux density than the random case due to lower applied field amplitude used to reach the desired flux density. The effect of temperature on the power loss at a given flux density can be calculated from the theory derived above. Susceptibility will in general decrease as , while the out-of-phase component will decrease further due to lower and . Thus higher field will be needed for higher temperature to reach desired flux density, but the power loss will decrease. Based on these calculations, the susceptibility and losses of perpendicular and parallel aligned spherical single-domain FeNi3 nanoparticles seem very promising for micro-inductor applications.
The direct effect of dipolar interaction on the susceptibility can be illustrated by use of equation (15). We find that for a hypothetical FeNi3 201 nm diameter two-particle system, an increase (/decrease) of susceptibility due to the two particles being oriented along a chain parallel (/perpendicular) to the applied field will be of the order of 0.2 (0.1) % of the total susceptibility at 30 vol% (assuming that the volume fraction is ). The effect due to the effective field therefore appears rather small, and the change holds true for particles with similar high moment. However, dipole fields of neighbouring particles might affect superparamagnetism [40] and thereby have a larger effect on susceptibility. Hence, we deem direct interaction effects on susceptibility to be of relatively low importance for materials with 30 vol% of particles, as long as the particles are still superparamagnetic.
The effect of dipolar interactions on the superparamagnetic relaxation is not yet fully described in the literature [40]. It has been suggested that dipolar interactions can slightly lower the effective anisotropy barrier in low volume fraction scenarios[41], while for dense aggregates it is often assumed to drastically increase effective anisotropy, consistently with the observed suppression of superparamagnetism [19, 42, 43, 40]. From energy considerations we argue that larger particle size with lower volume fraction are preferred, as the increase in uniaxial anisotropy energy is lower than the decrease in dipolar interaction energy, at a fixed nanocomposite susceptibility.
@@ IV Discussion
We have shown how to calculate the susceptibility for magnetic nanoparticle materials. Our theoretical framework provides novel results for particle alignment, cubic anisotropy, and effects from dipolar interactions that are not yet fully accounted for in the literature [44, 31]. Moreover, the dependency of particle susceptibility on the applied field direction with respect to the easy axis for uniaxial anisotropy case is interesting for application in power electronics magnetics, as it can be used to tune susceptibility and losses for different materials. We note that the susceptibility of cubic anisotropy particles cannot be tuned by alignment. However, as reported above (Fig. 4) most studies find a mild uniaxial anisotropy in spherical nanoparticles of cubic materials. Since magnetic properties, such as coercive field, are similar for the cubic and random uniaxial cases when not in the fully blocked state, it will challenging to distinguish cubic from uniaxial anisotropy when particles are not aligned [45, 26, 13].
@@
Magnetic nanoparticles have previously been investigated experimentally for use as inductor core materials in the MHz range [7, 8, 9, 10, 11, 12, 13, 14, 18, 16], (see Fig. 1). Our theoretical framework predicts susceptibility values in quantitative agreement with the general trends expounded by the many experimental reports available in the literature [12, 13, 8, 9]. This supports the predictive power of our model. The sparse mismatch observed between the model predictions and the data in some situations is ascribed to uncertainties in particle size distributions and material parameters (especially anisotropy and saturation magnetization, which may vary with size).
@@
The comparison between the experimental in-phase susceptibilities from [7, 8, 9, 10, 11, 12, 13, 14] (shown in Fig. 1) and our model highlights that most studies so far may have not optimized particle materials for the specific applications targeted. Hence, there is room for improvement, and the opportunity to break through the region of interest shown by the shaded area in Fig. 1. For example, some materials show little or no hysteresis [8, 9], which should be a sign of superparamagnetism, but these materials have only relatively low susceptibilities. This is connected to an insufficient saturation magnetisation of the particles (e.g. from oxides) and/or too dilute particle concentrations. For several other materials the nanoparticles were packed too densely or were aggregated [7, 10, 13, 14], which effectively renders the particles blocked and results in large coercive field of several thousand kA/m. Consequently, the resulting materials behave much like those comprising larger particles, i.e. with lower susceptibility and larger coercivity (see for example [11]). One reference reports a high particle susceptibility of 122, using relatively densely packed 8 nm Fe particles [12], which fits well with the results from Fig. 4 given that the resulting particle size is effectively larger than 8 nm. This is promising for micro-inductor applications.
@@
Nanocomposite materials thus show the potential to reach the area of interest of susceptibility above 50 in the frequency range above 2 MHz as indicated in Fig. 1. Based on our model, we propose that optimised particles in the form of spherical 9 nm Fe (or FeCo) or 20 nm FeNi3 particles, have the properties needed to develop inductor cores of the next generation. A difficulty in the applicability of our theoretical framework is that effective particle property values are somewhat sparse. Hence, the reported effective particle values in our evaluation of superparamagnetic particles may not be exact for all given sizes. Moreover, it is not always possible to synthesize particles in the desired size range, optimally monodisperse, and with the specific properties, concentration, or easy axis orientation. A relevant case study would be an uniaxial anisotropy particle accompanied by a small degree of shape anisotropy perpendicular to the uniaxial anisotropy, if they can be synthesized. This combination should produce a lower effective anisotropy barrier [46].
@@
The alignment of particles with the effective easy axis parallel to the applied field has been investigated experimentally [26, 13]. However, it is not clear whether the observed increase in susceptibility is primarily a result of alignment or of aggregation/chain formation. In fact, any form of clustering might block the particles or, at least, change significantly their dynamic behaviour [40]. To our knowledge, magnetic interactions in the derivation of susceptibilities typically are introduced as an effective field [31] resulting in susceptibility enhancement. On the other hand, Monte-Carlo simulations including the anisotropic dipolar interactions between nanoparticles indicated a susceptibility suppression for increasing volume fraction of particles [47]. Our theoretical framework sheds some light on the matter: it clarifies that for superparamagnetic particles the interactions and their effects depend on particle-assembly orientation with respect to the applied field, as shown in Eq. (15) and therefore depends on particle packing. However, the effect of interparticle interactions is only included in Eq. (15) and not included in the results presented in Figs. 3-6. The complex effects of dipolar interparticle interactions on susceptibility would be relevant to explore further in future studies.
@@
Dipolar interactions could potentially have also an effect in the frequency dependence because of the increased effective anisotropy. We note that many materials presented in Fig. 1 are relatively dense materials, and reported susceptibilities of low coercivity materials still seem to have stable susceptibility, although low, even at high frequencies [9, 12]. The applied field can change the superparamagnetic relaxation, allowing for faster response in applied field than without [48, 30]. This could explain the seemingly high blocking frequency seen for some densely packed materials [12, 9].
@@ V CONCLUSIONS
We have developed a comprehensive theoretical framework based on statistical mechanics to calculate nanoparticle susceptibility. The framework accounts for the effects of size, shape, anisotropy, saturation magnetization, and inter-particle interactions.
Our exploration of parameter space evidenced how nanoparticles with large saturation magnetisation and low anisotropy are the most suited for inductor cores in power electronics applications. Moreover, we determined that optimal particles should be as large as possible but remain superparamagnetic at the target operation frequency. For this, a narrow particle size distribution is preferable. In relation to particle shape, we found that elongation can increase susceptibility to some degree, but due to the increase in effective anisotropy, it is accompanied by an increase in out-of-phase component, which is associated to magnetic hysteresis losses. Therefore, spherical or near-spherical particle shapes would be most optimal for inductor cores. Losses per induced magnetisation are minimized for materials where the uniaxial anisotropy axes of the particles are aligned and not randomly distributed.
As an example of optimised particles, we show that a material with 30 vol% of spherical 201 nm FeNi3 particles could potentially have a susceptibility above 209 (487 for aligned particles) with lower losses than state-of-the-art ferrite magnetic core materials at 2 MHz operation.
The comparison between our model predictions and a broad experimental data set compiled from the literature, shows an excellent agreement, firmly supporting the quantitative predictive power of the model. Sub-optimal susceptibilities for high-frequency inductor core materials (´20-50) have been reported for nanocomposites in most experimental studies. This appears as a consequence of a poor choice of materials with low saturation magnetization, imperfect sizing of the particles, and either too dense or too dilute nanocomposites, or too dense aggregates with dipolar interactions suppressing the susceptibility. Our paper demonstrates that better nanocomposite materials are achievable to improve the performance of micro-fabricated inductor cores at high frequencies where bulk ferrite materials such as TDK’s PC200 cannot operate.
@@ Acknowledgements.
The authors thank Ron B. Goldfarb for stimulating discussions, and the Independent Research Fund Denmark for financial support (project HiFMag, grant number 9041-00231B).References
- Yawger [2022] R. Yawger, PSMA power technology roadmap, IEEE Power Electronics Magazine 9, 10 (2022).
@@
- Zambach et al. [2025] M. Zambach, M. Varón, M. R. Almind, M. Knaapila, Z. Ouyang, M. Beleggia, and C. Frandsen, Demagnetisation effects in single-domain particles (2025), arXiv:2502.03394 [cond-mat.mes-hall] .
- Stoner and Wohlfarth [1948] E. C. Stoner and E. P. Wohlfarth, A mechanism of magnetic hysteresis in heterogeneous alloys, Philosophical Transactions of the Royal Society of London Series A-mathematical and Physical Sciences 240, 599 (1948).
@@
- Walker et al. [1993] M. Walker, P. I. Mayo, K. OGrady, S. W. Charles, and R. W. Chantrell, The magnetic-properties of single-domain particles with cubic anisotropy .1. hysteresis loops, Journal of Physics-condensed Matter 5, 2779 (1993).
@@
- Ludwig et al. [2017] F. Ludwig, C. Balceris, and C. Johansson, The anisotropy of the ac susceptibility of immobilized magnetic nanoparticles-the influence of intra-potential-well contribution on the ac susceptibility spectrum, IEEE Transactions on Magnetics 53, 6101004 (2017).
- Raikher and Shliomis [1974] Y. Raikher and M. Shliomis, Theory of dispersion of the magnetic susceptibility of fine ferromagnetic particles, Sov. Phys. JETP 40, 526 (1974).
- Shliomis and Stepanov [1993] M. I. Shliomis and V. I. Stepanov, Frequency-dependence and long-time relaxation of the susceptibility of the magnetic fluids, Journal of Magnetism and Magnetic Materials 122, 176 (1993).
- Eisenstein and Aharoni [1977] I. Eisenstein and A. Aharoni, High-order superparamagnetic relaxation-times, Physica B and C 86-88, 1429 (1977).
- Carrey et al. [2011] J. Carrey, B. Mehdaoui, and M. Respaud, Simple models for dynamic hysteresis loop calculations of magnetic single-domain nanoparticles: Application to magnetic hyperthermia optimization, Journal of Applied Physics 109, 083921 (2011).
@@
- Bohra et al. [2023] M. Bohra, S. V. Battula, V. Alman, A. Annadi, and V. Singh, Design of various Ni–Cr nanostructures and deducing their magnetic anisotropy, Applied Nanoscience (Switzerland) 13, 3205 (2023).
@@
- Kumari et al. [2021] K. Kumari, A. Kumar, J. E. Lee, and B. H. Koo, Investigating the origin of exchange bias effect in ferromagnetic FeNi nanoparticles prepared via controlled synthesis, Applied Nanoscience 13, 613 (2021).
- Durhuus et al. [2025] F. L. Durhuus, T. H. Van Biljevelt Rix, M. A. Glod, M. Beleggia, and C. Frandsen, Néel relaxation of magnetic nanoparticle clusters, IEEE Magnetics Letters 16, 6100105 (5 pp.) (2025).